# Too Hot to Handle: Searching for Inflationary Particle Production in Planck Data

### v1

@@ I Introduction
Inflationary particle production is one such phenomenon that is present in a wide variety of multi-field scenarios, but has received less data analysis focus thus far than, e.g., primordial non-Gaussianity or primordial gravitational waves. The existence of multiple fields during inflation, which may not all be dynamical, is a generic possibility commonly predicted in UV models of particle physics, as well as string theory. A multi-field potential energy landscape involving all these fields can take various forms. Importantly, there are key observational differences that allow us to distinguish among different landscapes depending on the masses of fields. For example, light fields with masses much smaller than the inflationary Hubble scale $H_{I}$ can contribute to both curvature and isocurvature fluctuations. Particles with masses of order $H_{I}$ are produced less frequently but can give striking oscillatory signatures; this has been the focus of the recent “cosmological collider” program [3, 4]. On the other hand, extremely heavy fields are typically not excited, remaining stuck at the bottom of their potential, and can be “integrated out” from the inflationary dynamics.
@@
More specifically, we perform a matched-filter analysis using the Planck PR4 data [20]. We utilize the fact that searches for localized hot- or coldspots share analogies with existing searches for galaxy clusters in CMB maps via the thermal Sunyaev-Zel’dovich (tSZ) effect, i.e., inverse-Compton scattering of CMB photons off hot electrons in the intracluster medium [21]. [^1] State-of-the-art current tSZ cluster catalogs include Refs. [22, 23, 24]. In both cases, an angular profile for the signal is specified: in the hotspot case, the profile is computed from first-principles inflationary calculations [17, 18], whilst in the tSZ cluster case, the profile is modeled using functional forms calibrated by deep X-ray or tSZ observations of individual clusters [25, 26, 27]. Given the theoretical profile shape and knowledge of the data covariance, one can then construct and apply a matched filter [28] to the CMB maps to search for features with this profile. [^2] Analogous methods hold for point source detection, where the relevant profile is simply the instrument’s beam profile. A key difference between the hotspot and tSZ analyses, however, is that the inflationary hotspots possess the same blackbody spectral energy distribution (SED) as the (other) primary CMB fluctuations, whilst the tSZ effect produces a characteristic non-blackbody spectral distortion. One can thus exploit not only the angular profile of the signal, but also its SED, in such searches, via the use of multi-frequency matched-filter (MMF) techniques [29, 30]. Below, we adapt existing MMF tools that have been used for tSZ cluster-finding in Planck data to search for inflationary particle production hotspots, taking advantage of these analysis similarities.
@@ II Hotspots from Particle Production
$\begin{split}m_{\sigma}(\phi)\approx m_{\sigma}(\phi_{\star})+{1\over 2}m_{\sigma}^{\prime\prime}(\phi_{\star})(\phi-\phi_{\star})^{2}+\cdots,\end{split}$ (1)
@@
$\begin{split}{\cal L}_{\sigma}=-{1\over 2}(\partial_{\mu}\sigma)^{2}-{1\over 2}\left((g\phi-\mu)^{2}+M_{0}^{2}\right)\sigma^{2},\end{split}$ (2)
@@
$\begin{split}m_{\sigma}(\phi)\approx M_{0}+{g^{2}\over 2M_{0}}\dot{\phi}_{0}^{2}(t-t_{\star})^{2},\end{split}$ (3)
@@ III Methodology
In this work, we consider $10$ choices of $\eta_{\star}$ logarithmically spaced in $[10,1000]\,\,\mathrm{Mpc}$, as well as $10$ linearly spaced $\eta_{\rm HS}$ values, each satisfying $\chi_{\rm HS}\in[\chi_{\rm rec}-\eta_{\star},\chi_{\rm rec}+\eta_{\star}]$ and $0\leq\chi_{\rm HS}\leq\eta_{0}$, for comoving distance $\chi(\eta)\equiv\eta_{0}-\eta$. These conditions ensure that the particle production events lie in the observable Universe, and directly affect the last-scattering surface.
@@
We apply a number of cuts to limit contamination of primordial hotspots from late-time contributions. For Planck analyses, one could directly utilize public catalogs for this purpose [e.g., 39, 24]; to ensure consistency with the simulation-based tests below, we here measure them directly from the dataset (though we additionally compare any Planck detection candidates to the public catalogs in the below). We first find all point sources with $\mathrm{SNR}\geq 10$ in any of the six frequency channels, using a modified version of the above algorithm (analyzing each dataset in turn with a beam-convolved point-source template). The union of these catalogs defines a full-sky point-source map (here computed at $N_{\rm side}=4096$), which we use as a mask, excising regions within $3\sigma_{{\rm beam},\nu}$ via a recursive search, given the beam-size $\sigma_{{\rm beam},\nu}$ in channel $\nu$ (which increases with decreasing frequency, from roughly $2^{\prime}$ at 857 GHz to $4^{\prime}$ at 100 GHz). [^6] Recall that $\sigma_{\rm beam}={\rm FWHM}_{\rm beam}/\sqrt{8\ln 2}$, where FWHM is the full-width at half-maximum.
@@
Secondly, we perform a matched-filter analysis for tSZ clusters following [24], using the pipeline above, as implemented in szifi [40, 35, 34]. This uses a tSZ SED in the MMF of Eq. (6), and $15$ values of the cluster size, $\theta_{500}\in[5,50]^{\prime}$ (for the pressure profile of [25]). Following [34], we compute the SNR iteratively, masking out previously detected clusters when estimating the data covariance $\mathsf{C}_{\nu\nu^{\prime}}(l)$. Here, templates are applied for $l\in[100,2500]$ (matching previous works) and we apply an SNR threshold of $4.5$ (as in [33, 24]), additionally dropping any clusters within $5\sigma_{{\rm beam},\nu}$ of a point source. This catalog is used to post-process the hotspot catalog: we reject any candidates within $10^{\prime}$ of a tSZ cluster.
@@
Finally, we apply a conservative Galactic mask to the output catalog, removing any potential sources within the brightest $\approx 40\%$ of the sky in the highest frequency channel. This helps to avoid contamination from dusty clumps in the Milky Way and Magellanic Clouds, as described in the Planck Compton-$y$ map paper [41]. In concert with the above procedures, this leaves a total of $57.2\%$ of the sky for analysis (straightforwardly computed by injecting points and calculating the fraction not removed by masking operations). The end result of these procedures is a catalog of hotspot candidates, which can then be used for analysis (or for validation, if one injects a set of fake hotspots). An additional check (which we consider in the Planck section below) is to repeat the analysis using a standard (non-multi-frequency) matched filter applied to a component-separated CMB temperature map, rather than applying an MMF to the frequency maps; assuming an effective component separation method, this should have comparable SNR, since the hotspots have a blackbody SED.
@@ IV Validation Tests
We now proceed to test the above methodology using synthetic data. For this purpose, we use a single Planck NPIPE simulation [20], [^7] Available on NERSC:
@@ IV.1 Single Hotspots
Figure 2: Validation tests of the hotspot-finding pipeline, as applied to a single NPIPE simulation contaminated with $300$ isotropic hotspots across ten values of the size parameter $\eta_{\star}$, with distance $\eta_{\rm HS}$ fixed to $\eta_{\rm rec}$. Left panel: fraction of injected sources recovered at $\gtrsim 5\sigma$ as a function of the coupling parameter $g$. This asymptotes to $\approx 60\%$ due to masking, reflecting the fraction of sources obscured by late-time effects and Galactic contamination. Relative to the masked input catalog, the completeness is $100\%$ for sufficiently large $g,\eta_{\star}$. Larger hotspots are easier to detect: we are limited on small scales by the Planck beam, and on large scales by the flat-sky assumption and the Galactic mask. Middle panel: estimated values of $g$ from the injected sources. Though the uncertainty is large for small $\eta_{\star}$ (where hotspots are difficult to detect, with no sources for $\eta_{\star}<25\,\,\mathrm{Mpc}$), we find generally unbiased results at larger $\eta_{\star}$. Right panel: estimated values of $\eta_{\star}$ from the injected sources. As before, these are uncertain for small sources, but converge for large $g,\eta_{\star}$. We add small displacements along the abscissa for clarity.
@@
We begin by creating a map of $300$ hotspots at unit $g$, isotropically distributed on the full sky. Via Eq. (II), hotspots are defined on an $N_{\rm side}=2048$ HEALPix map, convolved with the Planck beam as before. For simplicity, we fix $\eta_{\rm HS}=\eta_{\rm rec}$ (i.e., place hotspots on the last-scattering surface [cf., 17]) but use $10$ values of $\eta_{\star}$ (each with $30$ hotspots). To avoid confusion in the inference, we ensure that all hotspots are separated by at least $1\,\,\mathrm{deg}$, such that they can be treated independently. This map is then combined linearly with the NPIPE temperature map at each frequency for $g\in[10,20,30,50]$ (following initial testing), and analyzed according to the above pipeline, after extraction of the point source and tSZ catalogs. Using $10$ $\eta_{\star}$ choices, the matched-filter analysis requires $\approx 20$ CPU-hours for each value of $g$, and can be trivially parallelized.
@@
In Fig. 2, we show the main results of the simulated single-hotspot analysis: the properties of injected hotspots recovered by our pipeline as a function of their (true) size, $\eta_{\star}^{\rm true}$. [^8] We declare a hotspot to be “recovered” if there is a $\geq 5\sigma$ candidate whose center lies within $10^{\prime}$ of the true center. If there are multiple (which commonly occurs around noisy hotspots), we take parameters from the candidate with largest SNR. From the first plot, we observe that our pipeline achieves good completeness for sufficiently large $g$ and $\eta_{\star}$; relative to the unmasked catalog of inputs, we find a recovery rate of $(64\pm 7)\%$ for $\eta_{\star}>25\,\,\mathrm{Mpc}$ and $g>20$, with the remaining hotspots hidden by the Galactic mask and secondary sources, such as tSZ clusters. If one instead defines the completeness with respect to the masked catalog, this increases to $100\%$, such that we recover all the inputs. As expected, hotspots with larger $g$ are easier to detect, with the completeness saturating at $g\approx 20$ for $\eta_{\star}>100\,\mathrm{Mpc}$. As $\eta_{\star}$ decreases, the completeness reduces significantly; small sources yield small temperature perturbations [cf., 17] and are hard to find given the Planck noise and beam. Hotspots with $\eta_{\star}>1000\,\,\mathrm{Mpc}$ (and angular sizes $\gtrsim 5\,\,\mathrm{deg}$) may also be detectable, though these are difficult to analyze since one cannot assume the flat-sky limit, may suffer from significant Galactic-plane contamination, and primary CMB cosmic variance is large on these scales.
@@ IV.2 Pairwise Hotspots
Next, we generate a map of pairwise hotspots, following a similar procedure to [18], though generalized outside the flat-sky limit. We first draw a single hotspot position $(\hat{\mathbf{n}}_{\rm HS,1},\chi_{\rm HS,1})$ given the aforementioned bounds, using a uniform distribution in Cartesian space. A second point is then drawn from a ball of radius $\eta_{\star}$ about the first (uniformly in Cartesian coordinates), and the pair is rejected if $\chi_{\rm HS,2}$ is not within $\eta_{\star}$ of the recombination surface (or outside the observable Universe). This procedure is repeated to generate $300$ pairs of points with known $\eta_{\rm HS,1,2}$ (which is discretized to match the values used in the analysis), including $10$ values of $\eta_{\star}$. These are used to construct a full-sky template map from Eq. (II), which is co-added to the NPIPE simulation, as before. The resulting map is then analyzed via the above pipeline. Importantly, the synthetic dataset contains hotspot pairs whilst the matched filter assumes single hotspots: this approach tests whether our method is applicable in the theory-motivated scenario of a distribution of pairs.
@@
We comment that in the context of the model in Eq. (2), values of $g\gg\sqrt{4\pi}$ are non-perturbatively large. However, in scenarios with more than one field during inflation, an “effective” value of $g\sim 10-100$ can be realized without violating such perturbativity bounds [42] whilst maintaining the same phenomenological hotspot properties. Therefore, to be model-agnostic we consider $g\sim 10-100$ here.

### v2

@@ I Introduction
Inflationary particle production is one such phenomenon that is present in a wide variety of multi-field scenarios, but has received less data analysis focus thus far than, e.g., primordial non-Gaussianity or primordial gravitational waves. The existence of multiple fields during inflation, which may not all be dynamical, is a generic possibility commonly predicted in UV models of particle physics, as well as string theory. A multi-field potential energy landscape involving all these fields can take various forms. Importantly, there are key observational differences that allow us to distinguish among different landscapes depending on the masses of fields. For example, light fields with masses much smaller than the inflationary Hubble scale $H_{I}$ can contribute to both curvature and isocurvature fluctuations. Particles with masses of order $H_{I}$ are produced less frequently but can give striking oscillatory signatures; this has been the focus of the recent “cosmological collider” program [3, 4, 5]. On the other hand, extremely heavy fields are typically not excited, remaining stuck at the bottom of their potential, and can be “integrated out” from the inflationary dynamics.
@@
More specifically, we perform a matched-filter analysis using the Planck PR4 data [21]. We utilize the fact that searches for localized hot- or coldspots share analogies with existing searches for galaxy clusters in CMB maps via the thermal Sunyaev-Zel’dovich (tSZ) effect, i.e., inverse-Compton scattering of CMB photons off hot electrons in the intracluster medium [22]. [^1] State-of-the-art current tSZ cluster catalogs include Refs. [23, 24, 25]. In both cases, an angular profile for the signal is specified: in the hotspot case, the profile is computed from first-principles inflationary calculations [18, 19], whilst in the tSZ cluster case, the profile is modeled using functional forms calibrated by deep X-ray or tSZ observations of individual clusters [26, 27, 28]. Given the theoretical profile shape and knowledge of the data covariance, one can then construct and apply a matched filter [29] to the CMB maps to search for features with this profile. [^2] Analogous methods hold for point source detection, where the relevant profile is simply the instrument’s beam profile. A key difference between the hotspot and tSZ analyses, however, is that the inflationary hotspots possess the same blackbody spectral energy distribution (SED) as the (other) primary CMB fluctuations, whilst the tSZ effect produces a characteristic non-blackbody spectral distortion. One can thus exploit not only the angular profile of the signal, but also its SED, in such searches, via the use of multi-frequency matched-filter (MMF) techniques [30, 31]; this additionally allows the hotspots and thermal SZ clusters to be practically distinguished. Below, we adapt existing MMF tools that have been used for tSZ cluster-finding in Planck data to search for inflationary particle production hotspots, taking advantage of these analysis similarities.
@@ II Hotspots from Particle Production
$\begin{split}m_{\sigma}(\phi)\approx m_{\sigma}(\phi_{\star})+\frac{1}{2}m_{\sigma}^{\prime\prime}(\phi_{\star})(\phi-\phi_{\star})^{2}+\cdots,\end{split}$ (1)
@@
$\begin{split}{\cal L}_{\sigma}=-\frac{1}{2}(\partial_{\mu}\sigma)^{2}-\frac{1}{2}\left((g\phi-\mu)^{2}+M_{0}^{2}\right)\sigma^{2},\end{split}$ (2)
@@
$\begin{split}m_{\sigma}(\phi)\approx M_{0}+\frac{g^{2}}{2M_{0}}\dot{\phi}_{0}^{2}(t-t_{\star})^{2},\end{split}$ (3)
@@ III Methodology
In this work, we consider $10$ choices of $\eta_{\star}$ logarithmically spaced in $[10,1000]\,\,\mathrm{Mpc}$, as well as $10$ linearly spaced $\eta_{\rm HS}$ values, each satisfying $\chi_{\rm HS}\in[\chi_{\rm rec}-\eta_{\star},\chi_{\rm rec}+\eta_{\star}]$ and $0\leq\chi_{\rm HS}\leq\eta_{0}$, for comoving distance $\chi(\eta)\equiv\eta_{0}-\eta$. The conditions on $\eta_{\rm HS}$ ensure that the particle production events lie in the observable Universe, and directly affect the last-scattering surface. Furthermore, the lower limit on $\eta_{\star}$ is set by the Planck beam: hotspots with $\eta_{\star}\lesssim 10\,\mathrm{Mpc}$ cannot be distinguished from point sources, except through their frequency dependence.
@@
We apply a number of cuts to limit contamination of primordial hotspots from late-time contributions. For Planck analyses, one could directly utilize public catalogs for this purpose [e.g., 40, 25]; to ensure consistency with the simulation-based tests below, we here measure them directly from the dataset (though we additionally compare any Planck detection candidates to the public catalogs in the below). We first find all point sources with $\mathrm{SNR}\geq 10$ in any of the six frequency channels, using a modified version of the above algorithm (analyzing each dataset in turn with a beam-convolved point source template). The union of these catalogs defines a full-sky point source map (here computed at $N_{\rm side}=4096$), which we use as a mask, excising regions within $3\sigma_{{\rm beam},\nu}$ via a recursive search, given the beam-size $\sigma_{{\rm beam},\nu}$ in channel $\nu$ (which increases with decreasing frequency, from roughly $2^{\prime}$ at 857 GHz to $4^{\prime}$ at 100 GHz). [^6] Recall that $\sigma_{\rm beam}={\rm FWHM}_{\rm beam}/\sqrt{8\ln 2}$, where FWHM is the full-width at half-maximum. Whilst most Planck analyses exclude all point sources with $\mathrm{SNR}\geq 5$, we use a less conservative choice in this work matching the Planck tSZ cluster analysis [25]. This ensures that we do not ‘over-mask’ and miss any hotspot candidates; to check for residual contamination, we compare the hotspot catalog to the set of $\mathrm{SNR}\geq 5$ point sources in the post-processing step.
@@
Secondly, we perform a matched-filter analysis for tSZ clusters following [25], using the pipeline above, as implemented in szifi [41, 36, 35]. This uses a tSZ SED in the MMF of Eq. (6), and $15$ values of the cluster size, $\theta_{500}\in[5,50]^{\prime}$ (for the pressure profile of [26]). Following [35], we compute the SNR iteratively, masking out previously detected clusters when estimating the data covariance $\mathsf{C}_{\nu\nu^{\prime}}(l)$. Here, templates are applied for $l\in[100,2500]$ (matching previous works) and we apply an SNR threshold of $4.5$ (as in [34, 25]), additionally dropping any clusters within $5\sigma_{{\rm beam},\nu}$ of a point source. This catalog is used to post-process the hotspot catalog by rejecting any candidates within $10^{\prime}$ of a tSZ cluster. This approach is, by construction, conservative: we consider only hotspot candidates that are not flagged by the point source or tSZ searches, rather than attempting to distinguish between the various types of source.
@@
Finally, we apply a conservative Galactic mask to the output catalog, removing any potential sources within the brightest $\approx 40\%$ of the sky in the highest frequency channel. This helps to avoid contamination from dusty clumps in the Milky Way and Magellanic Clouds, as described in the Planck Compton-$y$ map paper [42]. In concert with the above procedures, this leaves a total of $57.2\%$ of the sky for analysis (straightforwardly computed by injecting points and calculating the fraction not removed by masking operations). The end result of these procedures is a catalog of hotspot candidates, which can then be used for analysis (or for validation, if one injects a set of fake hotspots). An additional check (which we consider in the Planck section below) is to repeat the analysis using a standard (non-multi-frequency) matched filter applied to a component-separated CMB temperature map, rather than applying an MMF to the frequency maps; assuming an effective component separation method, this further reduces contamination from tSZ clusters and point sources, but should yield comparable SNR, since the hotspots have a blackbody SED.
@@ IV Validation Tests
We now proceed to test the above methodology using synthetic data. For this purpose, we use a single Planck FFP10/npipe simulation [21], [^7] Available on NERSC:
@@ IV.1 Single Hotspots
Figure 2: Validation tests of the hotspot-finding pipeline, as applied to a single FFP10/npipe simulation contaminated with $300$ isotropic hotspots across ten values of the size parameter $\eta_{\star}$, with distance $\eta_{\rm HS}$ fixed to $\eta_{\rm rec}$. Left panel: fraction of injected sources recovered at $\gtrsim 5\sigma$ as a function of the coupling parameter $g$. This asymptotes to $\approx 60\%$ due to masking, reflecting the fraction of sources obscured by late-time effects and Galactic contamination. Relative to the masked input catalog, the completeness is $100\%$ for sufficiently large $g,\eta_{\star}$. Larger hotspots are easier to detect: we are limited on small scales by the Planck beam, and on large scales by the flat-sky assumption and the Galactic mask. Middle panel: estimated values of $g$ from the injected sources. Though the uncertainty is large for small $\eta_{\star}$ (where hotspots are difficult to detect, with no sources for $\eta_{\star}<25\,\,\mathrm{Mpc}$), we find generally unbiased results at larger $\eta_{\star}$. Right panel: estimated values of $\eta_{\star}$ from the injected sources. As before, these are uncertain for small sources, but converge for large $g,\eta_{\star}$. We add small displacements along the abscissa for clarity.
@@
We begin by creating a map of $300$ hotspots at unit $g$, isotropically distributed on the full sky. Via Eq. (II), hotspots are defined on an $N_{\rm side}=2048$ HEALPix map, convolved with the Planck beam as before. For simplicity, we fix $\eta_{\rm HS}=\eta_{\rm rec}$ (i.e., place hotspots on the last-scattering surface [cf., 18]) but use $10$ values of $\eta_{\star}$ (each with $30$ hotspots). To avoid confusion in the inference, we ensure that all hotspots are separated by at least $1\,\,\mathrm{deg}$, such that they can be treated independently. This map is then combined linearly with the FFP10/npipe temperature map at each frequency for $g\in[10,20,30,50]$ (following initial testing). Given an analysis mask (built from the uninjected map to avoid the hotspots saturating the point source algorithm and thus being subtracted), we search for hotspots, using the matched-filter pipeline: using $10$ $\eta_{\star}$ choices, this requires $\approx 20$ CPU-hours for each value of $g$, and can be trivially parallelized.
@@
In Fig. 2, we show the main results of the simulated single-hotspot analysis: the properties of injected hotspots recovered by our pipeline as a function of their (true) size, $\eta_{\star}^{\rm true}$. [^8] We declare a hotspot to be “recovered” if there is a $\geq 5\sigma$ candidate whose center lies within $10^{\prime}$ of the true center. If there are multiple (which commonly occurs around noisy hotspots), we take parameters from the candidate with largest SNR. From the first plot, we observe that our pipeline achieves good completeness for sufficiently large $g$ and $\eta_{\star}$; relative to the unmasked catalog of inputs, we find a recovery rate of $(64\pm 7)\%$ for $\eta_{\star}>25\,\,\mathrm{Mpc}$ and $g>20$, with the remaining hotspots hidden by the analysis mask (which removes the Galactic plane and known secondary sources, such as tSZ clusters). If one instead defines the completeness with respect to the masked catalog, this increases to $100\%$, such that we recover all the inputs. As expected, hotspots with larger $g$ are easier to detect, with the completeness saturating at $g\approx 20$ for $\eta_{\star}>100\,\mathrm{Mpc}$. As $\eta_{\star}$ decreases, the completeness reduces significantly; small sources yield small temperature perturbations [cf., 18] and are swamped by the Planck detector noise; moreover, they would be indistinguishable from point sources, except through their frequency dependence. Hotspots with $\eta_{\star}>1000\,\,\mathrm{Mpc}$ (and angular sizes $\gtrsim 5\,\,\mathrm{deg}$) may also be detectable, though these are difficult to analyze since one cannot assume the flat-sky limit, may suffer from significant Galactic-plane contamination, and primary CMB cosmic variance is large on these scales.
@@ IV.2 Pairwise Hotspots
Next, we generate a map of pairwise hotspots, following a similar procedure to [19], though generalized outside the flat-sky limit. We first draw a single hotspot position $(\hat{\mathbf{n}}_{\rm HS,1},\chi_{\rm HS,1})$ given the aforementioned bounds, using a uniform distribution in Cartesian space. A second point is then drawn from a ball of radius $\eta_{\star}$ about the first (uniformly in Cartesian coordinates), and the pair is rejected if $\chi_{\rm HS,2}$ is not within $\eta_{\star}$ of the recombination surface (or outside the observable Universe). This procedure is repeated to generate $300$ pairs of points with known $\eta_{\rm HS,1,2}$ (which is discretized to match the values used in the analysis), including $10$ values of $\eta_{\star}$. These are used to construct a full-sky template map from Eq. (II), which is co-added to the FFP10/npipe simulation, as before. The resulting map is then analyzed via the above pipeline. Importantly, the synthetic dataset contains hotspot pairs whilst the matched filter assumes single hotspots: this approach tests whether our method is applicable in the theory-motivated scenario of a distribution of pairs.
@@
We comment that in the context of the model in Eq. (2), values of $g\gg\sqrt{4\pi}$ are non-perturbatively large. However, in scenarios with an extradimensional model of inflation, an “effective” value of $g\sim 10$ can be realized without violating such perturbativity bounds [43] whilst maintaining the same phenomenological hotspot properties. Scenarios involving tachyonic Higgs production [44] and quantum diffusion effects [45] may also lead to similar signatures. Therefore, to be model-agnostic we consider $g\sim 10-100$ here.